OpenStudy (anonymous):

what is the slope of x^2/16+y^2/36=1 at point (x,y)?

myininaya (myininaya):

It seems you are asked to find the "slope" formula aka the derivative of y w.r.t x

myininaya (myininaya):

so start by differentiating both sides

OpenStudy (anonymous):

I could find the slope with a point, but the point (x,y) is throwing my off.

myininaya (myininaya):

Do you know how to find the derivative?

myininaya (myininaya):

of x^2 and y^2 and a constant w.r.t. x?

myininaya (myininaya):

I'm asking if you know how to evaluate the following: \[\frac{d(x^2)}{dx} =? \\ \frac{d(y^2)}{dx} =? \\ \frac{d(1)}{dx}=?\]

OpenStudy (gorv):

welcome to OS @teach14 :)

OpenStudy (gorv):

just follow what he helped u and tell him/her ..so that he can help u further